OpenStudy (anonymous):

You get five packages of paper for $3.00. How much are you paying per package? Give your answer to the nearest cent.

OpenStudy (johnweldon1993):

So if you get 5 packages for 3 dollars....for 1 you pay ...x \[\frac{ 5 }{ 3 } = \frac{ 1 }{ x }\] cross multiply 5x = 3 and divide both sides by 5 x = 3/5 what is 3/5 in decimal form? .60...so 60 cents per package Hope that helps?

OpenStudy (anonymous):

yup thanks

OpenStudy (johnweldon1993):

no problem!